# Less is More: DocString Compression in Code Generation

### v2

@@ Abstract.
The widespread use of Large Language Models (LLMs) in software engineering has intensified the need for improved model and resource efficiency. In particular, for neural code generation, LLMs are used to translate function/method signature and DocString to executable code. DocStrings which capture user requirements for the code and used as the prompt for LLMs, often contains redundant information. Recent advancements in prompt compression have shown promising results in Natural Language Processing (NLP), but their applicability to code generation remains uncertain. Our empirical study show that the state-of-the-art prompt compression methods achieve only about 10% reduction, as further reductions would cause significant performance degradation. In our study, we propose a novel compression method, ShortenDoc, dedicated to DocString compression for code generation. Our extensive experiments on six code generation datasets, five open-source LLMs (1B to 10B parameters), and one closed-source LLM GPT-4o confirm that ShortenDoc achieves 25$\sim$40compression while preserving the quality of generated code, outperforming other baseline methods at similar compression levels. The benefit of this research is to improve efficiency and reduce the cost while maintaining the quality of the generated code, especially when calling third-party APIs, and is able to reduce the token processing cost by 25$\sim$40%.
@@ 1. Introduction
Large Language Models (LLMs) have emerged as a key tool in today’s software development, owning to their powerful language understanding and generation capabilities (Nam et al., 2024; Rasnayaka et al., 2024). In particular, they have shown great potential in a variety of SE tasks such as code generation (Liu et al., 2024), automated testing (Chen et al., 2024) and documentation (Dvivedi et al., 2024). However, with the wide deployment of LLMs, it has become a pressing challenge to improve the inference efficiency and to reduce the ever increasing demand of computational resources (Niu et al., 2024; Shi et al., 2024; Yang et al., 2024a).
@@ 12 $>$ add_numbers(2, 3)
DocString standing for ”documentation string” is a critical component that can significantly enhance the readability, maintainability and usability of the code. It plays a pivotal role in the code generation process, typically positioned at the onset of a function or method. DocStrings serve to articulate the code’s functionality, parameters, return values and potential exceptions that may arise (Dainese et al., 2024). Listing 1 presents an example to illustrate how DocString (Lines 2-14) elucidates the purpose of the function, its parameters, the type of return value it yields, the exceptions it might trigger, and offers a practical example showing how to invoke the function. These texts not only facilitate an intuitive grasp of the code for developers but also act as a direct reflection of user requirements.
In the realm of code generation leveraging LLMs, DocStrings are typically used as prompts. They guide the model to generate code that aligns with the specified requirements and expectations (Miceli-Barone and Sennrich, 2017). However, existing DocString varies significantly in quality (Poudel et al., 2024). Some DocString may be overly lengthy and contain a large amount of redundancy, which may prevent the model from accurately understanding user requirements. Although these redundant docstring do not exceed the length of the maximum context that LLM can now handle, they may also increase the computational cost of the model training and inference, thus reduce their efficiency. In addition, when users generate code by invoking APIs from third-party LLMs, shorter prompt can greatly reduce their financial cost. Therefore, optimizing the quality of DocString is key for improving the efficiency of code generation and reducing costs.
@@
(1) Compression efficacy: these compression methods show limited effectiveness in compressing DocStrings. When we attempted to increase the compression rate above 10%, there was a significant decrease in the quality of the generated code. We posit that this is largely because they fail to extract the code-related semantic information in DocStrings, leading to significant information loss.
(2) Inflexibility: existing methods require manual setting of the compression ratio, making it difficult to adapt to various code generation scenarios. Developers must adjust compression parameters based on specific situations, which requires specialized knowledge and is time-consuming. Moreover, it is hard to find the optimal balance of compression ratio and model efficacy (specifically Pass@1).
@@
  • We empirically show that the feasibility and limitations of DocString compression for code generation tasks.
@@ 3.1. Experiment Setup
Datasets. We experiment on six diverse and widely adopted datasets which can simulate real-world code generation tasks.
@@
Figure 1. The length distribution of DocString in these datasets.
@@
Existing Prompt Compression Methods. We explore two state-of-the-art prompt compression methods: Selective_Context (Li et al., 2023b) and LLMLingua2 (Pan et al., 2024). Selective_Context employs a base language model (such as GPT-2) to compute the self-information of lexical units such as sentences, phrases or tokens, which is then used to evaluate the informativeness. LLMLingua2 is a task-agnostic prompt compression technique, i.e., it does not depend on the information entropy of the prompt. It utilizes data distillation to create an effective prompt compression dataset and trains a classification model as a way to select tokens that can be compressed.
Evaluation Metrics. For code generation Pass@1 and Compression Ratio are considered. Pass@1 indicates the percentage of tasks for which the model produces a code snippet that successfully passes all the associated test cases on the first try. This metric is a direct reflection of the model’s accuracy and reliability in code generation scenarios. Compression Ratio quantifies the efficiency of the compression technique. It is defined as the percentage reduction in the length of the prompt after compression relative to the original length, i.e., $1-\frac{\|D^{\prime}\|}{\|D\|}$. A higher compression ratio indicates that more of the original prompt has been removed, which can lead to improvements in inference speed and reduced computational costs.
@@ 3.2. Empirical Findings
Table 1. Efficacy Comparison of Existing Compression Methods during Different Ratios
@@
In Table 1, we show the efficacy comparison of different LLMs after applying Selective_Context and LLMLingua2 compression techniques. In particular, for each model on different datasets, it shows the efficacy with the original prompts (0% compression) as well as the change in efficacy with increasing compression (10%, 20%, 30%, 40%).
(1) Effectiveness. As can be seen from the table, at lower compression rates, such as 10%, the efficacy degradation of most models is not significant. Taking the efficacy of CodeGeeX4 model on HumanEval dataset as an example, the original efficacy is 60.37%, the efficacy even can up to 62.20% at 10% compression rate using Selective_Context. Moreover, the efficacy can up to 62.80% at 10% compression rate using LLMLingua2. This shows that after compressing away some of the redundant information, the model still understands the task requirements better and generates correct code. This verifies that there is indeed redundant information in DocString that can be compressed without affecting the correctness of the generated code too much.
(2) Limitations. As the compression rate increases further, the efficacy degradation trend becomes obvious. For example, the efficacy of the CodeGeeX4 model on the HumanEval dataset decreases to 49.39% at a compression rate of 40% using Selective_Context, and to 37.20% at a compression rate of 40% using LLMLingua2. This suggests that existing compression methods, while removing more information, may also remove semantic information that is critical for the model to generate correct code.
@@
Table 2. The top-10 most frequent tokens in each dataset
@@
Compressed Token Analysis. In addition to analyzing the efficacy impact of prompt compression methods, we examine the specific tokens that are being removed. By reviewing the most frequent tokens that are compressed away, we can gain insights into the types of information that are considered redundant or less critical by the compression algorithms.
To this end, we extract the top-10 tokens with the highest frequency of occurrence for the two methods on the 10% compression ratio on the six datasets, respectively, and the results are shown in Table 2. We can find that the comparison between Selective_Context and LLMLingua2 reveals differences in the tokens each method targets for compression. This variation could be due to the distinct algorithms and criteria each method uses to evaluate the importance of tokens.
@@ 4. Approach
Building upon insights from the empirical research, we propose a novel compression technique ShortenDoc. Our approach consists of three main stages: Token Importance Sort, Search Space Formation and Constraint Specification.
[…]
@@ 4.1. Pre-processing
Inspired by the previous study (Wang et al., 2023b), we posit that the removal of line breaks and tabs from DocString does not significantly affect the efficacy of the code generation model. (An analysis will be discussed in Section 6.)
Moreover, upon examining the tokens that frequently undergo compression, as indicated in Table 2, we observe that the tokens compressed by LLMLingua2 exhibit regular patterns (such as ’the’, ’a’), and the model’s efficacy remains largely intact when compression is applied at a rate of 10%. This insight suggests that the compression of these tokens does not adversely affect model’s efficacy, leading us to categorize them as stop-words.
@@ 4.3. Search Space Construction
After ranking the tokens by their importance scores, we proceed to construct the search space $\mathcal{S}$ which comprises candidate tokens (sequences) that may be removed. To efficiently explore the search space and avoid suboptimal solutions due to local optima, we employ a Top-N strategy, i.e., to select the top N least important tokens $[d_{i_{1}},d_{i_{2}},\ldots,d_{i_{N}}]$. Our empirical observations in Table 1 indicate that, in certain scenarios, a higher compression rate may yield enhanced efficacy. Taking the DS-6.7B model as an example, on the CodeHarmony dataset, the efficacy of LLMLingua2 under 20% compression outperforms that under 10% compression. We hypothesize that the simultaneous compression of consecutive tokens could provide unexpected benefits over compressing individual tokens. As a result, for each $1\leq k\leq N$ we consider $k$-gram, i.e.,
@@ 4.4. Constraint Definition
Here, $\tau$ is the hyper-parameter and we also set 0.999 in our study.
@@ 5. Evaluation
5.1. RQ1: What is the effectiveness of our proposed ShortenDoc compared to existing prompt compression methods? (Efficacy Comparison)
We aim to evaluate the effectiveness of ShortenDoc in reducing DocString length without compromising the efficacy of code generation models. To ensure the robustness of our findings, we also include a closed-source state-of-the-art model GPT-4o in our analysis. Additionally, we introduce the Random method as a baseline. The focus of our analysis is on the efficacy metrics and the compression ratio across a spectrum of datasets and models, as presented in Table 3.
Table 3. Efficacy Comparison of Different Methods Across Various Datasets and Models
DataSet Method Ratio DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B GPT-4o
@@
Our results demonstrate that ShortenDoc consistently outperforms alternative methods across all datasets and models, particularly at higher compression rates. For instance, on the HumanEval dataset, ShortenDoc achieves the highest efficacy across all models, effectively maintaining or even improving code generation quality despite a 30% reduction in DocString length. This indicates that ShortenDoc can sustain, and in some cases enhance, model efficacy while significantly compressing input prompts.
Moreover, in several cases, ShortenDoc surpasses the uncompressed baseline. For instance, on the HumanEval dataset, DeepSeekCoder-6.7B, CodeQwen, and CodeGeeX4 can even achieve a higher Pass@1 value when compressed by ShortenDoc. This surprising result suggests that ShortenDoc not only preserves essential information but may also eliminate redundant or distracting content from the DocStrings, thereby improving model’s focus on relevant information during code generation.
@@
Furthermore, the fact that ShortenDoc maintains high efficacy across models of varying sizes and architectures indicates its generalizability. It effectively aids both open-source and closed-source models, including state-of-the-art systems like GPT-4o. This broad applicability enhances the practical value of ShortenDoc in real-world code generation applications.
Efficiency: FLOPs (Floating Point Operations) refers to the number of floating-point operations, an indicator used to measure the complexity of a model, reflecting the total number of floating-point operations required during its execution. FLOPs is closely related to the input length and structure of the model. The purpose of docstring compression is to reduce computational costs and enhance model efficiency. Therefore, we utilize the FLOPs metric to evaluate the computational load of different models before and after compression across various datasets.
@@
Summary of RQ1 ShortenDoc consistently outperforms existing compression methods across multiple datasets and models, effectively reducing DocString length without compromising, and sometimes even enhancing code generation efficacy.
@@
DataSet Method Ratio DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
Considering that the datasets chosen in RQ1 are all in Python, to explore the generalizability of ShortenDoc, we select four other programming languages, i.e., C++, Go, Java, JavaScript in the HumanEval-X (Zheng et al., 2023) dataset for our experiments. Since these four datasets share DocString with HumanEval, we directly migrate the compressed DocString obtained for HumanEval in RQ1 to these datasets. The detailed experimental results are shown in Table 5.
The results reveal that ShortenDoc maintains its superior performance across all four programming languages when compared to baseline methods (e.g., Random compression, SelectiveContext, and LLMLingua2). Despite the structural and syntactical differences among these languages, ShortenDoc consistently achieves higher performance metrics, demonstrating its generalization ability to retain the critical information in DocStrings.
@@
At the same time, we observe a slight degradation in model performance when directly migrating the compressed DocStrings obtained in RQ1 (Python) to other programming languages such as C++, Go, Java, and JavaScript. While ShortenDoc consistently outperformed baseline methods, its performance did not always fully match the results achieved with the original uncompressed DocStrings in these new languages.
This performance degradation can be attributed to the fact that the base language model used in our approach, CodeGPT, is specifically trained and optimized for Python. The model’s understanding of the syntax, structure, and semantics of Python-based code influences how DocStrings are compressed. When these compressed DocStrings are applied to other programming languages, certain language-specific nuances may not be as effectively preserved, leading to a small drop in code generation performance.
@@
Summary of RQ2 While ShortenDoc generalizes well across programming languages, a slight performance degradation is observed when migrating compressed DocStrings from Python to other languages. This can be attributed to the base model, CodeGPT, being trained specifically for Python, which may not fully capture the nuances of other languages. Nonetheless, the generalization capability of ShortenDoc was thoroughly demonstrated by its performance across C++, Go, Java, and JavaScript in the HumanEval-X dataset. By directly migrating compressed DocStrings from Python datasets, ShortenDoc consistently outperformed baseline methods, proving its adaptability and robustness across multiple programming languages.
Figure 3. Case Study of Different DocString Compression Methods

### v3

@@ Abstract.
The widespread use of Large Language Models (LLMs) in software engineering has intensified the need for improved model and resource efficiency. In particular, for neural code generation, LLMs are used to translate function/method signature and DocString to executable code. DocStrings, which capture user requirements for the code and are typically used as the prompt for LLMs, often contain redundant information. Recent advancements in prompt compression have shown promising results in Natural Language Processing (NLP), but their applicability to code generation remains uncertain. Our empirical study show that the state-of-the-art prompt compression methods achieve only about 10% reduction, as further reductions would cause significant performance degradation. In our study, we propose a novel compression method, ShortenDoc, dedicated to DocString compression for code generation. Our experiments on six code generation datasets, five open-source LLMs (1B to 10B parameters) and one closed-source LLM GPT-4o confirm that ShortenDoc achieves 25–40% compression while preserving the quality of generated code, outperforming other baseline methods at similar compression levels. The benefit of this method is to improve efficiency and reduce the token processing cost while maintaining the quality of the generated code, especially when calling third-party APIs.
@@ 1. Introduction
Large Language Models (LLMs) have emerged as a key tool in today’s software development, owning to their powerful language understanding and generation capabilities (Nam et al., 2024; Rasnayaka et al., 2024). In particular, they have shown great potential in a variety of SE tasks such as code generation (Liu et al., 2024), automated testing (Chen et al., 2024) and documentation (Dvivedi et al., 2024). However, with the wide deployment of LLMs, it has become a pressing challenge to improve the inference efficiency and to reduce the ever increasing demand of computational resources (Niu et al., 2024; Shi et al., 2024; Yang et al., 2024a). This challenge manifests in two aspects: (1) the computational overhead in terms of FLOPs during model inference, and (2) the financial cost associated with API calls to third-party LLMs for code generation.
@@ 12 $>$ add_numbers(2, 3)
DocString (documentation string), a string literal embedded in source code to document specific features or functionalities, serves as a critical component in software development. It significantly enhances code readability, maintainability and usability. Typically positioned at the beginning of a function or method, DocString plays a pivotal role in the code generation process by articulating the code’s functionality, parameters, return values and potential exceptions that may arise (Dainese et al., 2024). Listing 1 presents an example to illustrate how DocString (Lines 2-14) elucidates the purpose of the function, its parameters, the type of return value it yields, the exceptions it might trigger, and offers a practical example showing how to invoke the function.
Motivation. In the realm of code generation leveraging LLMs, DocStrings are typically used as prompts. They guide the model to generate code that aligns with the document specific features or functionalities (Miceli-Barone and Sennrich, 2017). However, existing DocString varies significantly in quality (Poudel et al., 2024). Some DocString may be overly lengthy and contain a large amount of redundancy, which may prevent the model from accurately understanding user requirements. Although these redundant DocString do not exceed the length of the maximum context that LLM can now handle, they may also increase the computational cost of the model inference, thus reduce their efficiency. For instance, the computational cost of the model inference is usually proportional to the length of the input prompt. The longer the DocString, the higher the computational cost of model inference. According to OpenAI founder Sam Altman, GPT-4o processes approximately 200 billion tokens daily [^1] https://x.com/sama/status/1815437745550172617, while Baidu CEO Robin Li revealed (at Baidu World 2024) that ERNIE model handles over 1.5 billion calls per day [^2] https://ir.baidu.com/news-releases/news-release-details/baidu-announces-third-quarter-2024-results/. At such massive processing scales, even a modest 25%-40% compression of user inputs could significantly reduce overall computational costs and energy consumption in the long run.
In addition, when users generate code by invoking APIs from third-party LLMs, shorter prompt can greatly reduce their financial cost. Therefore, optimizing the quality of DocString is one of the key factors for improving the efficiency of code generation and reducing costs. For instance, the costs of LLM API are associated with input tokens (i.e., prefilling) and output tokens (i.e., decoding), where input tokens are usually 1–5 times cheaper than output tokens (varying across different LLM providers). For example, Meta’s Llama 3.1 405B model charges at the same rate for both input and output tokens [^3] https://llama3-1.com/price/; OpenAI’s models price input tokens at 1/4 of output tokens [^4] https://openai.com/api/pricing/; Anthropic’s models set input tokens at 1/5 of output token costs [^5] https://www.anthropic.com/pricing. Assuming one usage consumes 512 input tokens and 512 output tokens on average, with a 50% input token reduction, the end-to-end cost saving is $(512\times 1\times 50\%)/(512\times 1+512\times 4)=10\%$ according to OpenAI’s pricing model, and 25% according to Meta’s pricing model. While the cost reduction for a single API call might seem modest due to the pricing difference between input and output tokens across LLM providers, the cumulative economic benefits become substantial in large-scale deployment scenarios or when synthesizing large-scale code generation datasets using LLMs.
Moreover, DocString compression can be complementary to other efficiency-oriented techniques in code generation, such as efficient decoding strategies (Sun et al., 2024a, b) and optimized programming language grammar (Sun et al., 2024c). Meanwhile, the insights gained from DocString compression in zero-shot code generation scenarios can also provide theoretical foundations for more complex retrieval-augmented generation (RAG) based code generation tasks.
@@
(1) Compression efficacy: these compression methods show limited effectiveness in compressing DocStrings. As evidenced by our experimental results in Section 3 (Table 1), increasing the compression rate beyond 10% may incur a significant decrease in the quality of the generated code. We posit that this is largely because they fail to extract the code-related semantic information in DocStrings, leading to significant information loss.
(2) Inflexibility: existing methods require manual setting of the compression ratio, making it difficult to adapt to various code generation scenarios. Developers must adjust compression parameters based on specific situations, which requires specialized knowledge and is time-consuming. Moreover, it is hard to find the optimal balance of compression ratio and model efficacy (specifically Pass@1, which measures the percentage of tasks for which the model produces a correct output on its first attempt).
@@
  • We empirically show the feasibility and limitations of DocString compression for code generation tasks.
@@
To facilitate reproducibility, all source code and experimental data are released at https://github.com/NTDXYG/ShortenDoc.
@@ 3.1. Experiment Setup
Datasets. We experiment on six diverse and widely adopted datasets which can simulate real-world code generation tasks. Our dataset selection follows two principles: (1) including dataset variants derived from the same base to evaluate robustness, and (2) covering different sources and complexity levels to ensure diversity.
For the first principle, we select HumanEval and its two variants from EvoEval (Subtle and Creative). While these datasets share similar base problems, they differ in how requirements are presented. This setup allows us to evaluate ShortenDoc’s robustness in handling different expressions of the same functional requirement.
For the second principle, we select datasets from varied sources and at different complexity levels: MBPP contains human-written entry-level problems, CodeHarmony provides LLM-synthesized high-quality samples, and BigCodeBench presents practical and challenging programming tasks. The diversity ensures a comprehensive evaluation across different scenarios. In particular,
@@
Figure 1. The length distribution of DocString in these datasets. Notice for each dataset, the percentage represents the proportion of samples falling within a specific length range relative to the total number of samples. The length range indicates the number of tokens in the DocString after tokenization.
@@
Existing Prompt Compression Methods. We explore two state-of-the-art prompt compression methods: Selective_Context (Li et al., 2023b) and LLMLingua2 (Pan et al., 2024).
Selective_Context employs a base language model (such as GPT-2) to compute the self-information of lexical units such as sentences, phrases or tokens, which is then used to evaluate the informativeness. Specifically, it operates at the noun phrase level by first using NLTK for lexical analysis, then applying GPT tokenizer for subword tokenization, and finally merging tokens into noun phrases using spacy, which explains why some compressed tokens appear as subwords (e.g., ”tu” for ”tuples”) in Fig. 3.
LLMLingua2 is a task-agnostic prompt compression technique, i.e., it does not depend on the information entropy of the prompt. It utilizes data distillation to create an effective prompt compression dataset and trains a classification model as a way to select tokens that can be compressed. Unlike Selective_Context, LLMLingua2 operates at the complete word level, using a BERT classifier to identify redundant tokens.
Evaluation Metrics. For code generation Pass@1 and Compression Ratio are considered. Pass@1 indicates the percentage of tasks for which the model produces a code snippet that successfully passes all the associated test cases on the first try. This metric is a direct reflection of the model’s accuracy and reliability in code generation scenarios. Compression Ratio quantifies the efficiency of the compression technique. It is defined as the percentage reduction in the length of the prompt after compression relative to the original length, i.e., $1-\frac{\|D^{\prime}\|}{\|D\|}$, where $D$ is the original DocString and $D^{\prime}$ is the compressed one. A higher compression ratio indicates that more of the original prompt has been removed, which can lead to improvements in inference speed and reduced computational costs.
@@ 3.2. Empirical Findings
Table 1. Pass@1 Comparison of Existing Compression Methods during Different Ratios
@@
In Table 1, we show the comparison of Pass@1 across different LLMs after applying Selective_Context and LLMLingua2 compression techniques. In particular, for each model on different datasets, it shows Pass@1 with the original prompts (0% compression) as well as the change in Pass@1 with increasing compression (10%, 20%, 30%, 40%).
(1) Effectiveness. As can be seen from the table, with a 10% compression rate, most models maintain comparable or even improved Pass@1 scores relative to their original performance. Taking the Pass@1 of CodeGeeX4 model on HumanEval dataset as an example, the original Pass@1 is 60.37%, the Pass@1 even can up to 62.20% at 10% compression rate using Selective_Context. Moreover, the Pass@1 can up to 62.80% at 10% compression rate using LLMLingua2.
However, this improvement is not universal across all datasets. Notably, on the BigCodeBench dataset, most models show a decrease in Pass@1 scores after compression. We attribute this phenomenon to the inherent complexity and diversity of DocStrings across different datasets.
These mixed results suggest that, while some DocStrings contain redundant information that can be safely compressed without compromising performance, the effectiveness of compression is highly dependent on the dataset characteristics. This observation highlights the importance of considering dataset-specific features when applying compression techniques.
(2) Limitations. As the compression rate increases further, the Pass@1 degradation trend becomes obvious. For example, the Pass@1 of the CodeGeeX4 model on the HumanEval dataset decreases to 49.39% at a compression rate of 40% using Selective_Context, and to 37.20% at a compression rate of 40% using LLMLingua2. This suggests that existing compression methods, while removing more information, may also remove semantic information that is critical for the model to generate correct code.
@@
Table 2. The top-10 most frequently removed tokens in each dataset
@@
Compressed Token Analysis. In addition to analyzing the compression ratio of prompt compression methods, we examine the specific tokens that are being removed. By reviewing the most frequent tokens that are compressed away, we can gain insights into the types of information that are considered redundant or less critical by the compression algorithms.
To this end, we extract the top-10 most frequently removed tokens with the highest frequency of occurrence for the two methods on the 10% compression ratio on the six datasets, respectively, and the results are shown in Table 2. We can find that the comparison between Selective_Context and LLMLingua2 reveals differences in the tokens each method targets for compression. This variation could be due to the distinct algorithms and criteria each method uses to evaluate the importance of tokens.
@@ 4. Approach
Building upon insights from the empirical research, we propose a novel compression technique ShortenDoc, which is shown in Fig. 2. Our approach consists of four main stages: Step 1: Pre-processing, Step 2: Token Importance Sort, Step 3: Search Space Construction, and Step 4: Constraint Definition.
Given a prompt consisting of a signature $T$ and a DocString $D$, our goal is to compress the DocString into $D^{\prime}$ while preserving the overall semantics and effectiveness in guiding LLMs for code generation.
@@
Specifically, in our approach, we opt for BPE (Byte Pair Encoding) tokenizer’s subword level, which aligns with the native tokenization of most modern LLMs while providing fine-grained compression control, and eliminates the need for additional token conversion steps, reducing computational overhead. Furthermore, subword tokenization offers finer-grained control over compression. For instance, when compressing technical terms or compound words in DocStrings, subword-level operations can preserve critical parts while removing less essential components.
@@ 4.1. Pre-processing
Inspired by the previous study (Wang et al., 2023b), we posit that the removal of line breaks and tabs from DocString does not significantly affect Pass@1 of the code generation model. (An analysis will be discussed in Section 6.)
Moreover, upon examining the tokens that frequently undergo compression, as indicated in Table 2, we observe that the tokens compressed by LLMLingua2 exhibit regular patterns (such as ’the’, ’a’), and the model’s Pass@1 score remains largely intact when compression is applied at a rate of 10%. This insight suggests that the compression of these tokens does not adversely affect model’s Pass@1 score, leading us to categorize them as stop-words.
@@ 4.3. Search Space Construction
After ranking the tokens by their importance scores, we proceed to construct the search space $\mathcal{S}$ which comprises candidate tokens (sequences) that may be removed. To efficiently explore the search space and avoid suboptimal solutions due to local optima, we employ a Top-N strategy, i.e., to select the top N least important tokens $[d_{i_{1}},d_{i_{2}},\ldots,d_{i_{N}}]$. Our empirical observations in Table 1 indicate that, in certain scenarios, a higher compression rate may yield enhanced Pass@1 score. Taking the DS-6.7B model as an example, on the CodeHarmony dataset, the Pass@1 of LLMLingua2 under 20% compression outperforms that under 10% compression. We hypothesize that the simultaneous compression of consecutive tokens could provide unexpected benefits over compressing individual tokens. As a result, for each $1\leq k\leq N$ we consider $k$-gram, i.e.,
@@ 4.4. Constraint Definition
Here, $\tau$ is the hyper-parameter which is set to be 0.999 in our study.
@@ 5. Evaluation
5.1. RQ1: What is the effectiveness of our proposed ShortenDoc compared to existing prompt compression methods? (Pass@1 Comparison)
We aim to evaluate the effectiveness of ShortenDoc in reducing DocString length without compromising the Pass@1 score of code generation models. To ensure the robustness of our findings, we also include a closed-source state-of-the-art model GPT-4o in our analysis. Additionally, we introduce the Random method as a baseline. The focus of our analysis is on the Pass@1 score and the compression ratio across a spectrum of datasets and models, as presented in Table 3.
The compression ratios for different datasets are carefully determined through preliminary experiments, considering the following factors:
  • Dataset Characteristics: Different datasets exhibit varying DocString patterns and information density. For instance, MBPP’s DocStrings contain more redundant information, allowing a higher compression ratio (38%), while CodeHarmony’s more concise DocStrings necessitate a lower ratio (25%).
  • Fair Comparison: To ensure a fair comparison, we apply the same compression ratio to all methods within each dataset. While different methods might achieve their optimal performance at different ratios, using consistent ratios within datasets allows a direct comparison of compression effectiveness.
Table 3. Pass@1 Comparison of Different Methods Across Various Datasets and Models
DataSet Method Ratio LLMs
DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B GPT-4o
@@
Our results demonstrate that ShortenDoc consistently outperforms alternative methods across all datasets and models, particularly at higher compression rates. For instance, on the HumanEval dataset, ShortenDoc achieves the highest Pass@1 score across all models, effectively maintaining or even improving code generation quality despite a 30% reduction in DocString length. This indicates that ShortenDoc can sustain, and in some cases enhance, model Pass@1 score while significantly compressing input prompts.
Moreover, in several cases, ShortenDoc surpasses the uncompressed baseline. For instance, on the HumanEval dataset, DeepSeekCoder-6.7B, CodeQwen and CodeGeeX4 can even achieve a higher Pass@1 score when compressed by ShortenDoc. This surprising result suggests that ShortenDoc not only preserves essential information but may also eliminate redundant or distracting content from the DocStrings, thereby improving model’s focus on relevant information during code generation.
@@
Furthermore, the fact that ShortenDoc maintains a high Pass@1 score across models of varying sizes and architectures indicates its generalizability. It effectively aids both open-source and closed-source models, including state-of-the-art systems like GPT-4o. This broad applicability enhances the practical value of ShortenDoc in real-world code generation applications.
Efficiency: To comprehensively evaluate the efficiency of ShortenDoc, we analyze its performance from three perspectives:
(1) GPU Memory Usage: GPU memory consumption is a critical resource constraint in deploying LLMs for code generation tasks. To evaluate the memory efficiency of our approach, we compare the GPU memory usage between using compressed and uncompressed DocStrings.
It is worth noting that, while our method employs CodeGPT-py-adapted as the base model for compression, its memory footprint is negligible compared to the deployment of LLMs. Specifically, CodeGPT-py-adapted has only 124M parameters and requires merely 0.23GB of GPU memory under the torch.bfloat16 precision. This overhead is insignificant when compared to the substantial memory requirements of modern LLMs used for code generation.
(2) FLOPs Analysis: FLOPs (Floating Point Operations) refers to the number of floating-point operations, an indicator used to measure the complexity of a model, reflecting the total number of floating-point operations required during its execution. FLOPs is closely related to the input length and structure of the model. The purpose of DocString compression is to reduce computational costs and enhance model efficiency. Therefore, we utilize the FLOPs metric to evaluate the computational load of different models before and after compression across various datasets.
@@
(3) Inference Time: End-to-end inference time is a crucial metric for real-world applications, directly impacting user experience and system throughput. To assess the practical benefits of our compression approach, we conduct comprehensive timing measurements under controlled experimental conditions using NVIDIA RTX 3090 GPU. All models were configured to use the torch.bfloat16 precision with both input and output lengths set to 512 tokens.
The results show that the code generation time varies significantly across different model scales. The DeepSeekCoder model with 1.3B parameters requires approximately 9.21 seconds per generation, while larger models around 7B parameters take about 17.77 seconds. For comparison, when using the GPT-4 API, the generation time typically ranges from 5 to 10 seconds, though it may fluctuate depending on the server load.
The time overhead introduced by ShortenDoc’s compression process is relatively modest, averaging around 2 seconds per DocString. This additional preprocessing time proves to be a worthwhile investment. In practical applications, the total processing pipeline (compression plus inference) remains highly efficient.
Summary of RQ1 ShortenDoc consistently outperforms existing compression methods across multiple datasets and models, effectively reducing DocString length without compromising, and sometimes even enhancing code generation Pass@1 score.
@@
DataSet Method Ratio LLMs
DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
Considering that the datasets chosen in RQ1 are all in Python, to explore the generalizability of ShortenDoc, we select four other programming languages, i.e., C++, Go, Java, JavaScript in the HumanEval-X (Zheng et al., 2023) dataset for our experiments.
Our choice of HumanEval-X is primarily driven by the status quo. At present, most publicly available code generation datasets for non-Python languages are variants of HumanEval, as it has become a de facto benchmark in the field. While we acknowledge that using variations of a single dataset might introduce potential biases, HumanEval-X represents one of the few comprehensive, well-validated multilingual code generation benchmarks available to the research community.
Since these four datasets share DocString with HumanEval, we directly migrate the compressed DocString obtained for HumanEval in RQ1 to these datasets. The detailed experimental results are shown in Table 5.
The results reveal that ShortenDoc maintains its superior performance across all four programming languages when compared to baseline methods (e.g., Random compression, SelectiveContext and LLMLingua2). Despite the structural and syntactical differences among these languages, ShortenDoc consistently achieves higher performance metrics, demonstrating its generalization ability to retain the critical information in DocStrings.
@@
At the same time, we observe a slight degradation in Pass@1 scores when directly migrating the compressed DocStrings obtained in RQ1 (Python) to other programming languages such as C++, Go, Java and JavaScript. While ShortenDoc consistently outperformed baseline methods, its performance did not always fully match the results achieved with the original uncompressed DocStrings in these new languages.
This performance degradation can be attributed to the fact that the base language model used in our approach, CodeGPT, is specifically trained and optimized for Python. The model’s understanding of the syntax, structure and semantics of Python-based code influences how DocStrings are compressed. When these compressed DocStrings are applied to other programming languages, certain language-specific nuances may not be as effectively preserved, leading to a small drop in code generation performance.
@@
Summary of RQ2 While ShortenDoc generalizes well across programming languages, a slight performance degradation is observed when migrating compressed DocStrings from Python to other languages. This can be attributed to the base model, CodeGPT, being trained specifically for Python, which may not fully capture the nuances of other languages. Nonetheless, the generalization capability of ShortenDoc was thoroughly demonstrated by its performance across C++, Go, Java and JavaScript in the HumanEval-X dataset. By directly migrating compressed DocStrings from Python datasets, ShortenDoc consistently outperformed baseline methods, proving its adaptability and robustness across multiple programming languages.
Figure 3. Example of Different DocString Compression Methods